OpenStudy (anonymous):

here are three geometric series. a) 2+1+0.5+... b) 75+30+12+... c) 240-60+15-3.75+.... 1) For each of the series i) find the common ratio r ii) use GDC to calculate the values of S10, S15, S20. Write full values that you see on your GDC screen 2) Do you notice any patterns? why do you think this is happening 3) Now use your GDC to calculate the values of S50 for each series. Do you think your calculator is correct? Explain why or why not

OpenStudy (anonymous):

What is GDC?

OpenStudy (anonymous):

grafic calculator

Parth (parthkohli):

\[S(n) = a\frac{1-r^n}{1-r }\]

OpenStudy (anonymous):

so how do you answer the queston

OpenStudy (anonymous):

you can divide the second term by the first to find r. \[r=\frac{a_2}{a_1}\] So for the first one r = 1/2

Parth (parthkohli):

\[S(10) = a\frac{1 - r^{10}}{1-r }\]

Parth (parthkohli):

\[S(15) = a\frac{1-r^{15}}{1-r}\]

Parth (parthkohli):

\[S(20) = a\frac{1 - r^{20}}{1-r}\]

OpenStudy (anonymous):

ok thansks